find the center and radius of a circle, given the following equation of the circle
x^2+y^2=100

Answer by piyu(4)

if the equation of the circle is of the type x^2+y^2=r^2, then the circle has its centre at the origin and its radius is r
your equation was
x^2+y^2=100
here the radius is the sq. root of 100 (i.e.)10